# RegExplainer: Generating Explanations for Graph Neural Networks in Regression Tasks

### v3

@@ B.2 Property 2
where $\mathcal{L}_{\text{N}}$ denotes the NCE loss. By inserting the optimal $\text{sim}\left(Y^{*},Y\right)$ into Eq. (11), we could get:
@@ Appendix C Graph Mix-up Approach
Specifically, for a target graph $G_{a}$ in the original graph set to be explained, we generate the explanation sub-graph $G_{a}^{*}=G_{a}-G_{a}^{\Delta}$ from the explainer. To generate a graph in the same distribution of original $G_{a}$, we can randomly sample a graph $G_{b}$ from the original set, generate the explanation sub-graph of $G_{b}^{*}$ with the same explainer and retrieve its label-irrelevant graph $G_{b}^{\Delta}=G_{b}-G_{b}^{*}$. Then we could merge $G_{a}^{*}$ together with $G_{b}^{\Delta}$ and produce the mix-up explanation $G_{a}^{\text{(mix)}}$. Formally, we can have $G_{a}^{\text{(mix)}}=G_{a}^{*}+(G_{b}-G_{b}^{*})$.
Since we are using the edge weights mask to describe the explanation, we can denote $G_{a}$ and $G_{b}$ with the adjacency matrices ${\bm{A}}_{a}$ and ${\bm{A}}_{b}$, their edge weight mask matrices as ${\bm{M}}_{a}$ and ${\bm{M}}_{b}$. If $G_{a}$ and $G_{b}$ are aligned graphs with the same number of nodes, we could simply mix them up by ${\bm{M}}_{a}^{\text{(mix)}}={\bm{M}}_{a}^{*}+({\bm{I}}_{b}-{\bm{M}}_{b}^{*})$, where ${\bm{M}}$ denotes the weight of the adjacency matrix and ${\bm{I}}_{b}$ denotes the zero-ones matrix as weights of all edges in the adjacency matrix of $G_{b}$, where 1 represents the existing edge and 0 represents there is no edge between the node pair.
If $G_{a}$ and $G_{b}$ are not aligned with the same number of nodes, we can use a connection adjacency matrix ${\bm{A}}_{\text{conn}}$ and mask matrix ${\bm{M}}_{\text{conn}}$ to merge two graphs with different numbers of nodes. Specifically, the mix-up adjacency matrix could be formed as:
@@
And the mix-up mask matrix could be formed as:
@@
Finally, we could form $G_{a}^{\text{(mix)}}$ as $({\bm{X}}^{\text{(mix)}},{\bm{A}}_{a}^{\text{(mix)}}\odot{\bm{M}}_{a}^{\text{(mix)}})$, where ${\bm{X}}^{\text{(mix)}}=[{\bm{X}}_{a};{\bm{X}}_{b}]$. The detailed algorithm for mix-up is shown in Algorithm 1. In implantation, $\eta=|\mathcal{E}|*0.03$.

### v4

@@ B.2 Property 2
where $\mathcal{L}_{\text{N}}$ denotes the NCE loss. By inserting the optimal $\text{sim}\left(Y^{*},Y\right)$ into Eq. (11), we can get:
@@ Appendix C Graph Mix-up Approach
Specifically, for a target graph $G_{a}$ in the original graph set to be explained, we generate the explanation sub-graph $G_{a}^{*}=G_{a}-G_{a}^{\Delta}$ from the explainer. To generate a graph in the same distribution of original $G_{a}$, we can randomly sample a graph $G_{b}$ from the original set, generate the explanation sub-graph of $G_{b}^{*}$ with the same explainer and retrieve its label-irrelevant graph $G_{b}^{\Delta}=G_{b}-G_{b}^{*}$. Then we can merge $G_{a}^{*}$ together with $G_{b}^{\Delta}$ and produce the mix-up explanation $G_{a}^{\text{(mix)}}$. Formally, we can have $G_{a}^{\text{(mix)}}=G_{a}^{*}+(G_{b}-G_{b}^{*})$.
Since we are using the edge weights mask to describe the explanation, we can denote $G_{a}$ and $G_{b}$ with the adjacency matrices ${\bm{A}}_{a}$ and ${\bm{A}}_{b}$, their edge weight mask matrices as ${\bm{M}}_{a}$ and ${\bm{M}}_{b}$. If $G_{a}$ and $G_{b}$ are aligned graphs with the same number of nodes, we can simply mix them up by ${\bm{M}}_{a}^{\text{(mix)}}={\bm{M}}_{a}^{*}+({\bm{I}}_{b}-{\bm{M}}_{b}^{*})$, where ${\bm{M}}$ denotes the weight of the adjacency matrix and ${\bm{I}}_{b}$ denotes the zero-ones matrix as weights of all edges in the adjacency matrix of $G_{b}$, where 1 represents the existing edge and 0 represents there is no edge between the node pair.
If $G_{a}$ and $G_{b}$ are not aligned with the same number of nodes, we can use a connection adjacency matrix ${\bm{A}}_{\text{conn}}$ and mask matrix ${\bm{M}}_{\text{conn}}$ to merge two graphs with different numbers of nodes. Specifically, the mix-up adjacency matrix can be formed as:
@@
And the mix-up mask matrix can be formed as:
@@
Finally, we can form $G_{a}^{\text{(mix)}}$ as $({\bm{X}}^{\text{(mix)}},{\bm{A}}_{a}^{\text{(mix)}}\odot{\bm{M}}_{a}^{\text{(mix)}})$, where ${\bm{X}}^{\text{(mix)}}=[{\bm{X}}_{a};{\bm{X}}_{b}]$. The detailed algorithm for mix-up is shown in Algorithm 1. In implantation, $\eta=|\mathcal{E}|*0.03$.